OpenStudy (anonymous):

Find the equation of the tangent line to f (x) = x^2 -4 when X=5

OpenStudy (anonymous):

I found the slope of the tangent line to be = 2x.... so p in for x gives m=10..... but how to I find the y int to complete the equation for a kibe?

OpenStudy (anonymous):

A line**

OpenStudy (anonymous):

sustitude (x=5), in the given equation then solve for y

OpenStudy (anonymous):

Which equation do you substitute x=5 into to solve for the y int?

OpenStudy (anonymous):

sorry, network problem., in the original equation

OpenStudy (anonymous):

If I do that I get 21 and the answer is supposed to be -29

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

there is something wrong with my Pc, but b4 you find y int., you need to find the corresponding value of x=5, which is y=2, from there make a new equation, using the slope, and point(5,21)

OpenStudy (anonymous):

look, I have to go, the new equation will be \[y=10x-29\], then let x=0, to find y int.

OpenStudy (anonymous):

here's a better solution:\[f(5) = 5^2 - 4 = 21\] so your points are (5, 21). next find the slope fo the tangent line to the function X^2 - 4. Take the derivative of the function. The derivative is the slope of the tangent line. The slope is 2x. The plug in 5 to get 10. to get the equation of the tangent line, you have a point and a slope. so, use your point slope form: \[(y - y _{1}) = m(x-x _{1})\] so you have the following:\[y - 5 = 10(x-5)\] the rest is simple....just solve for y to get your equation to the tangent line.